# Discrete Probabilistic Inference as Control in Multi-path Environments

### v1

@@ 3 Bridging the gap between MaxEnt RL & GFlowNets
There exists a fundamental difference between MaxEnt RL and GFlowNets in the way the distributions induced by their (optimal) policies relate to the energy: in MaxEnt RL, this distribution is over how objects are being created (trajectories) as mentioned in Section 2.2, whereas a GFlowNet induces a distribution over the outcomes only (terminating states), the latter matching the requirements of the Gibbs distribution. In this section, we recall some existing connections between GFlowNets and MaxEnt RL (Tiapkin et al., 2024), and we establish new equivalences between existing MaxEnt RL algorithms and GFlowNet objectives.
@@ 3.1 Reward correction
To correct the bias illustrated in Figure 1 caused by multiple trajectories leading to the same terminating state, we can treat Equation 5 not as a sum but as an expectation over trajectories, by reweigthing $\pi^{*}(\tau)$ (which is constant for a fixed terminating state, by Equation 4 & Equation 2) with a probability distribution over these trajectories. Bengio et al. (2023) showed that such a distribution over complete trajectories can be defined by introducing a backward transition probability $P_{B}(s\mid s^{\prime})$, which is a distribution over the parents $s\in\mathrm{Pa}(s^{\prime})$ of any state $s^{\prime}\neq s_{0}$. Tiapkin et al. (2024) showed that the reward of the soft MDP can be modified based on $P_{B}$ in such a way that the corresponding optimal policy $\pi^{*}_{\mathrm{MaxEnt}}$ is equal to the GFlowNet policy $P_{F}$ in Equation 7. We restate and generalize this result in Theorem 3.1, where we show how this correction counteracts the effect of the marginalization in Equation 5, resulting in a terminating state distribution that matches the Gibbs distribution.
@@
where we used the convention $s_{T+1}=s_{f}$. Then the terminating state distribution associated with the optimal policy $\pi_{\mathrm{MaxEnt}}^{*}$ solution of Equation 3 satisfies $\pi^{*}(x)\propto\exp(-{\mathcal{E}}(x)/\alpha)$.
The proof of the theorem is available in Appendix A. Unlike Equation 2, the return now depends on the trajectory leading to $s_{T}$ via the second term in Equation 8. Interestingly, the temperature parameter $\alpha$ introduced in the MaxEnt RL literature (Haarnoja et al., 2017) finds a natural interpretation as the temperature of the Gibbs distribution. Note that the correction in Equation 8 only involves the backward probability of the whole trajectory $\tau$, making it also compatible even with non-Markovian $P_{B}$ (Shen et al., 2023; Bengio et al., 2023).
Tiapkin et al. (2024) only considered the case where the reward function of the soft MDP is sparse, and the correction with $P_{B}(s_{t}\mid s_{t+1})$ is added at each intermediate transition $s_{t}\rightarrow s_{t+1}$; we will go back to this setting in the following section. This correction of the reward is fully compatible with our observation in Section 2.2 that sampling terminating states with $\pi^{*}_{\mathrm{MaxEnt}}$ yields samples of Equation 1 when the soft MDP is a tree with the (uncorrected) reward in Equation 2, since in that case any state $s^{\prime}\neq s_{0}$ has a unique parent $s$, and thus $P_{B}(s\mid s^{\prime})=1$ as also observed by Tiapkin et al. (2024).
3.2 Equivalence between Path Consistency Learning & (Sub-)Trajectory Balance
@@
$r(s_{t},s_{t+1})$ $=\alpha\log P_{B}(s_{t}\mid s_{t+1})$ $r(s_{T},s_{f})$ $=-{\mathcal{E}}(s_{T}),$ (9)
as introduced in (Tiapkin et al., 2024). Theorem 3.1 suggests that solving the MaxEnt RL problem in Equation 3 with the corrected reward is comparable to finding a solution of a GFlowNet, as they both lead to a policy whose terminating state distribution is the Gibbs distribution. It turns out that there exists an equivalence between specific algorithms solving these two problems with our choice of reward function above: Path Consistency Learning (PCL; Nachum et al., 2017) for MaxEnt RL, and the Subtrajectory Balance objective (SubTB; Madan et al., 2023; Malkin et al., 2022) for GFlowNets. Both of these objectives operate at the level of partial trajectories of the form $\tau=(s_{m},s_{m+1},\ldots,s_{n})$, where $s_{m}$ and $s_{n}$ are not necessarily the initial and terminal states anymore.
@@
where $P_{B}$ is a backward transition probability, which we assume to be fixed here—although in general, $P_{B}$ may also be learned (Malkin et al., 2022). In addition to this objective, some boundary conditions on $F_{\phi}$ must also be enforced (depending on ${\mathcal{E}}$), similar to the ones introduced in Section 2.3; see Section B.1 for details. The following proposition establishes the equivalence between these two objectives, up to a normalization constant that only depends on the temperature $\alpha$, and provides a way to move from the policy/value function parametrization in MaxEnt RL to the policy/flow function parametrization in GFlowNets. Although similarities between these two methods have been mentioned in prior work (Malkin et al., 2022; Jiralerspong et al., 2024; Hu et al., 2024; Mohammadpour et al., 2024), we show here an exact equivalence between both objectives.
@@
The proof of this proposition is available in Section B.1. The equivalence between the value function in MaxEnt RL and the state flow function in GFlowNets was also found in (Tiapkin et al., 2024). When applied to complete trajectories, this also shows the connection between the Trajectory Balance objective (TB; Malkin et al., 2022), widely used in the GFlowNet literature, and PCL under our choice of corrected reward in Equation 9.
On the other end of the spectrum, if we apply this proposition to transitions in the soft MDP (i.e., subtrajectories of length 1), then we can obtain a similar equivalence between the Detailed Balance objective in GFlowNets (DB; Bengio et al., 2023), and the Soft Q-Learning algorithm (SQL; Haarnoja et al., 2017), via the Unified PCL perspective of Nachum et al. (2017) that uses a soft Q-function in order to simultaneously parametrize both the policy and the value function in Equation 10; see Corollary B.1 for a detailed statement. We note that this connection between DB & SQL was also mentioned in prior work (Tiapkin et al., 2024; Mohammadpour et al., 2024). A full summary of the connections resulting from Proposition 3.2 between different MaxEnt RL and GFlowNet objectives is available in Figure 2, with further details in Figure 6.
[…]
Figure 2: Equivalence between the objectives of MaxEnt RL algorithms, with corrected rewards defined in Equation 9, and the objectives in GFlowNets.
@@ 3.3 Soft Q-Learning with policy parametrization
In this section, we consider the case where all the states of the soft MDP are valid elements of the sample space ${\mathcal{X}}$ (in other words, all the states are terminating). Since every state is now associated with some energy, we can reshape the rewards (Ng et al., 1999) while still satisfying Equation 8 as
$r(s_{t},s_{t+1})$ $={\mathcal{E}}(s_{t})-{\mathcal{E}}(s_{t+1})+\alpha\log P_{B}(s_{t}\mid s_{t+1})$ $r(s_{T},s_{f})$ $=0,$ (13)
if we assume, without loss of generality, that ${\mathcal{E}}(s_{0})=0$ (any offset added to the energy function leaves Equation 1 unchanged). This is a novel setting that differs from Tiapkin et al. (2024), and was made possible thanks to our general statement in Theorem 3.1. We show in Proposition B.2 that with our choice of rewards above, in particular the fact that no reward is received upon termination, we can express the objective of Soft Q-Learning as a function of a policy $\pi_{\theta}$ parametrized by $\theta$, instead of a Q-function; we call this $\pi$-SQL. The objective can be written as ${\mathcal{L}}_{\pi\textrm{-}\mathrm{SQL}}(\theta)=\frac{1}{2}\mathbb{E}_{\pi_{b}}[\Delta_{\pi\textrm{-}\mathrm{SQL}}^{2}(s,s^{\prime};\theta)]$, where $\pi_{b}$ is an arbitrary distribution over transitions $s\rightarrow s^{\prime}$ such that $s^{\prime}\neq s_{f}$, and
@@
$\Delta_{\mathrm{M}\textrm{-}\mathrm{DB}}(s,s^{\prime};\theta)=\log\frac{\exp(-{\mathcal{E}}(s^{\prime})/\alpha)P_{B}(s\mid s^{\prime})P_{F}^{\theta}(s_{f}\mid s)}{\exp(-{\mathcal{E}}(s)/\alpha)P_{F}^{\theta}(s^{\prime}\mid s)P_{F}^{\theta}(s_{f}\mid s^{\prime})}.$ (15)
@@ 4 Related Work
Unlike standard reinforcement learning where an optimal policy may be completely deterministic (at least in the fully observable case; Sutton and Barto, 2018), MaxEnt RL seeks a stochastic policy that balances between reward maximization and maximal entropy of future actions (Ziebart, 2010; Fox et al., 2016). This type of entropy regularization falls into the broader domain of regularized MDPs (Geist et al., 2019). This can be particularly beneficial for improving exploration (Haarnoja et al., 2017) and for robust control under model misspecification (Eysenbach and Levine, 2022). Popular MaxEnt RL methods include Soft Q-Learning, Path Consistency Learning (Nachum et al., 2017), and Soft Actor-Critic (Haarnoja et al., 2018a) studied in this paper.
@@
Bengio et al. (2021) took inspiration from reinforcement learning and introduced GFlowNets as a solution for finding diverse molecules binding to a target protein. Since then, they have found applications in a number of domains in scientific discovery (Jain et al., 2022, 2023a; Mila AI4Science et al., 2023), leveraging diversity in conjunction with active learning (Jain et al., 2023b; Hernandez-Garcia et al., 2023), but also in combinatorial optimization (Zhang et al., 2023a, b), causal discovery (Deleu et al., 2022, 2023; Atanackovic et al., 2023), and probabilistic inference in general (Zhang et al., 2022b; Hu et al., 2023, 2024; Falet et al., 2024). Although they were framed differently, GFlowNets are also deeply connected to the literature on variational inference (Malkin et al., 2023; Zimmermann et al., 2022).
A number of works have recently alluded to connections between GFlowNets and (maximum entropy) RL. The only work to formally establish the connection is Tiapkin et al. (2024), which showed how the reward in MaxEnt RL can be corrected based on some backward transition probability to be equivalent to GFlowNets. Although their analysis is limited to the case where the reward function in the original soft MDP is sparse (i.e., the reward is only obtained at the end of the trajectory), they were the first to propose a correction applied at each intermediate transition as in Section 3.2. We generalized this in Theorem 3.1 with a correction at the level of the trajectories, which offers more flexibility in how the correction is distributed along the trajectory and allows intermediate rewards. Tiapkin et al. (2024) also showed similarities between GFlowNet objectives and MaxEnt RL algorithms, namely between Detailed Balance & Dueling Soft Q-Learning (Wang et al., 2016), and between Trajectory Balance & Policy Gradient (Schulman et al., 2017). The correspondence between TB and Policy Gradient was further expanded in (Anonymous, 2023), as a direct consequence of the connections between GFlowNets, variational inference (Malkin et al., 2023), and reinforcement learning (Weber et al., 2015). Finally, Mohammadpour et al. (2024) also introduced a correction that depends on $n(s)$ the number of (partial) trajectories to a certain state $s$, which can be learned by solving a second MaxEnt RL problem Equation 3 on an “inverse” MDP. This correction corresponds to a particular choice of backward transition probability $P_{B}(s_{t}\mid s_{t+1})=n(s_{t})/n(s_{t+1})$ in Equation 8, which has the remarkable property of maximizing the flow entropy.
@@ 5 Experimental results
We verify empirically the equivalences established in Section 3 on three domains: the inference over discrete factor graphs (Buesing et al., 2020), Bayesian structure learning of Bayesian networks (Deleu et al., 2022), and the generation of parsimonious phylogenetic trees (Zhou et al., 2024). In addition to Detailed Balance (and possibly its modified version) and Trajectory Balance on the one hand (GFlowNets), and Soft Q-Learning (possibly parametrized by a policy; see Section 3.3) and Path Consistency Learning on the other hand (MaxEnt RL), we also consider a discrete version of Soft Actor-Critic (Christodoulou, 2019), which has no natural conterpart in the GFlowNet literature. For all MaxEnt RL methods, we adjust the MDP to include the correction of the reward. Note that in all the domains considered here, intermediate rewards are available in the original MDP, meaning in particular that all instances of DB actually use the Forward-Looking formulation (Pan et al., 2023a). Additional experimental details are available in App. C.

### v2

@@ 3 Bridging the gap between MaxEnt RL & GFlowNets
There exists a fundamental difference between MaxEnt RL and GFlowNets in the way the distributions induced by their (optimal) policies relate to the energy: in MaxEnt RL, this distribution is over how objects are being created (trajectories) as mentioned in Section 2.2, whereas a GFlowNet induces a distribution over the outcomes only (terminating states), the latter matching the requirements of the Gibbs distribution. In this section, we recall some existing connections between GFlowNets and MaxEnt RL [Tiapkin et al., 2024], and we establish new equivalences between existing MaxEnt RL algorithms and GFlowNet objectives.
@@ 3.1 Reward correction
To correct the bias illustrated in Figure 1 caused by multiple trajectories leading to the same terminating state, we can treat Equation 5 not as a sum but as an expectation over trajectories, by reweigthing $\pi^{*}(\tau)$ (which is constant for a fixed terminating state, by LABEL:eq:distribution-trajectories & Equation 2) with a probability distribution over these trajectories. Bengio et al. [2023] showed that such a distribution over complete trajectories can be defined by introducing a backward transition probability $P_{B}(s\mid s^{\prime})$, which is a distribution over the parents $s\in\mathrm{Pa}(s^{\prime})$ of any state $s^{\prime}\neq s_{0}$. Tiapkin et al. [2024] showed that the reward of the soft MDP can be modified based on $P_{B}$ in such a way that the corresponding optimal policy $\pi^{*}_{\mathrm{MaxEnt}}$ is equal to the GFlowNet policy $P_{F}$ in Equation 7. We restate and generalize this result in Theorem 3.1, where we show how this correction counteracts the effect of the marginalization in Equation 5, resulting in a terminating state distribution that matches the Gibbs distribution.
@@
where we used the convention $s_{T+1}=s_{f}$. Then the terminating state distribution associated with the optimal policy $\pi_{\mathrm{MaxEnt}}^{*}$ solution of LABEL:eq:maxent-rl-problem satisfies $\pi^{*}(x)\propto\exp(-{\mathcal{E}}(x)/\alpha)$.
The proof of the theorem is available in Appendix A. Unlike Equation 2, the return now depends on the trajectory leading to $s_{T}$ via the second term in Equation 8. Interestingly, the temperature parameter $\alpha$ introduced in the MaxEnt RL literature [Haarnoja et al., 2017] finds a natural interpretation as the temperature of the Gibbs distribution. Note that the correction in Equation 8 only involves the backward probability of the whole trajectory $\tau$, making it also compatible even with non-Markovian $P_{B}$ [Shen et al., 2023, Bengio et al., 2023].
Tiapkin et al. [2024] only considered the case where the reward function of the soft MDP is sparse, and the correction with $P_{B}(s_{t}\mid s_{t+1})$ is added at each intermediate transition $s_{t}\rightarrow s_{t+1}$; we will go back to this setting in the following section. This correction of the reward is fully compatible with our observation in Section 2.2 that sampling terminating states with $\pi^{*}_{\mathrm{MaxEnt}}$ yields samples of Equation 1 when the soft MDP is a tree with the (uncorrected) reward in Equation 2, since in that case any state $s^{\prime}\neq s_{0}$ has a unique parent $s$, and thus $P_{B}(s\mid s^{\prime})=1$ as also observed by Tiapkin et al. [2024].
3.2 Equivalence between PCL & (Sub)TB
@@
$r(s_{t},s_{t+1})$ $=\alpha\log P_{B}(s_{t}\mid s_{t+1})$ (9)
$r(s_{T},s_{f})$ $=-{\mathcal{E}}(s_{T}),$
as introduced in [Tiapkin et al., 2024]. Theorem 3.1 suggests that solving the MaxEnt RL problem in LABEL:eq:maxent-rl-problem with the corrected reward is comparable to finding a solution of a GFlowNet, as they both lead to a policy whose terminating state distribution is the Gibbs distribution. It turns out that there exists an equivalence between specific algorithms solving these two problems with our choice of reward function above: Path Consistency Learning (PCL; Nachum et al., 2017) for MaxEnt RL, and the Subtrajectory Balance objective (SubTB; Madan et al., 2023, Malkin et al., 2022) for GFlowNets. Both of these objectives operate at the level of partial trajectories of the form $\tau=(s_{m},s_{m+1},\ldots,s_{n})$, where $s_{m}$ and $s_{n}$ are not necessarily the initial and terminal states anymore.
@@
$\Delta_{\mathrm{PCL}}(\tau;$ $\theta,\phi)=-V_{\mathrm{soft}}^{\phi}(s_{m})+V_{\mathrm{soft}}^{\phi}(s_{n})$ (10)
$+\sum_{t=m}^{n-1}\big{(}r(s_{t},s_{t+1})-\alpha\log\pi_{\theta}(s_{t+1}\mid s_{t})\big{)}.$
@@
$\Delta_{\mathrm{SubTB}}(\tau;\theta,\phi)=\log\frac{F_{\phi}(s_{n})\prod_{t=m}^{n-1}P_{B}(s_{t}\mid s_{t+1})}{F_{\phi}(s_{m})\prod_{t=m}^{n-1}P_{F}^{\theta}(s_{t+1}\mid s_{t})},$
where $P_{B}$ is a backward transition probability, which we assume to be fixed here—although in general, $P_{B}$ may also be learned [Malkin et al., 2022]. In addition to this objective, some boundary conditions on $F_{\phi}$ must also be enforced (depending on ${\mathcal{E}}$), similar to the ones introduced in Section 2.3; see Section B.1 for details. The following proposition establishes the equivalence between these two objectives, up to a normalization constant that only depends on the temperature $\alpha$, and provides a way to move from the policy/value function parametrization in MaxEnt RL to the policy/flow function parametrization in GFlowNets. Although similarities between these two methods have been mentioned in prior work [Malkin et al., 2022, Jiralerspong et al., 2024, Hu et al., 2024, Mohammadpour et al., 2024], we show here an exact equivalence between both objectives.
@@
The proof of this proposition is available in Section B.1. The equivalence between the value function in MaxEnt RL and the state flow function in GFlowNets was also found in [Tiapkin et al., 2024]. When applied to complete trajectories, this also shows the connection between the Trajectory Balance objective (TB; Malkin et al., 2022), widely used in the GFlowNet literature, and PCL under our choice of corrected reward in Equation 9.
On the other end of the spectrum, if we apply this proposition to transitions in the soft MDP (i.e., subtrajectories of length 1), then we can obtain a similar equivalence between the Detailed Balance objective in GFlowNets (DB; Bengio et al., 2023), and the Soft Q-Learning algorithm (SQL; Haarnoja et al., 2017), via the Unified PCL perspective of Nachum et al. [2017] that uses a soft Q-function in order to simultaneously parametrize both the policy and the value function in Equation 10; see Corollary B.1 for a detailed statement. We note that this connection between DB & SQL was also mentioned in prior work [Tiapkin et al., 2024, Mohammadpour et al., 2024]. A full summary of the connections resulting from Proposition 3.2 between different MaxEnt RL and GFlowNet objectives is available in Figure 2, with further details in Figure 6.
@@ 3.3 Soft Q-Learning with policy parametrization
In this section, we consider the case where all the states of the soft MDP are valid elements of the sample space ${\mathcal{X}}$ (in other words, all the states are terminating). Since every state is now associated with some energy, we can reshape the rewards [Ng et al., 1999] while still satisfying Equation 8 as
$r(s_{t},s_{t+1})$ $={\mathcal{E}}(s_{t})-{\mathcal{E}}(s_{t+1})+\alpha\log P_{B}(s_{t}\mid s_{t+1})$
$r(s_{T},s_{f})$ $=0,$ (12)
if we assume, without loss of generality, that ${\mathcal{E}}(s_{0})=0$ (any offset added to the energy function leaves Equation 1 unchanged). This is a novel setting that differs from Tiapkin et al. [2024], and was made possible thanks to our general statement in Theorem 3.1. We show in Proposition B.2 that with our choice of rewards above, in particular the fact that no reward is received upon termination, we can express the objective of Soft Q-Learning as a function of a policy $\pi_{\theta}$ parametrized by $\theta$, instead of a Q-function; we call this $\pi$-SQL. The objective can be written as ${\mathcal{L}}_{\pi\textrm{-}\mathrm{SQL}}(\theta)=\frac{1}{2}\mathbb{E}_{\pi_{b}}[\Delta_{\pi\textrm{-}\mathrm{SQL}}^{2}(s,s^{\prime};\theta)]$, where $\pi_{b}$ is an arbitrary distribution over transitions $s\rightarrow s^{\prime}$ such that $s^{\prime}\neq s_{f}$, and
$\Delta_{\pi\textrm{-}\mathrm{SQL}}(s,s^{\prime};\theta)=\alpha\big{[}$ $\log\pi_{\theta}(s^{\prime}\mid s)-\log\pi_{\theta}(s_{f}\mid s)$
$+\log\pi_{\theta}(s_{f}\mid s^{\prime})\big{]}-r(s,s^{\prime}).$ (13)
@@
$\Delta_{\mathrm{M}\textrm{-}\mathrm{DB}}(s,s^{\prime};\theta)$ (14)
$\qquad=\log\frac{\exp(-{\mathcal{E}}(s^{\prime})/\alpha)P_{B}(s\mid s^{\prime})P_{F}^{\theta}(s_{f}\mid s)}{\exp(-{\mathcal{E}}(s)/\alpha)P_{F}^{\theta}(s^{\prime}\mid s)P_{F}^{\theta}(s_{f}\mid s^{\prime})}.$
@@ 4 Related Work
Unlike standard reinforcement learning where an optimal policy may be completely deterministic (at least in the fully observable case; Sutton and Barto, 2018), MaxEnt RL seeks a stochastic policy that balances between reward maximization and maximal entropy of future actions [Ziebart, 2010, Fox et al., 2016]. This type of entropy regularization falls into the broader domain of regularized MDPs [Geist et al., 2019]. This can be particularly beneficial for improving exploration [Haarnoja et al., 2017] and for robust control under model misspecification [Eysenbach and Levine, 2022]. Popular MaxEnt RL methods include Soft Q-Learning, Path Consistency Learning [Nachum et al., 2017], and Soft Actor-Critic [Haarnoja et al., 2018a] studied in this paper.
@@
Bengio et al. [2021] took inspiration from reinforcement learning and introduced GFlowNets as a solution for finding diverse molecules binding to a target protein. Since then, they have found applications in a number of domains in scientific discovery [Jain et al., 2022, 2023a, Mila AI4Science et al., 2023], leveraging diversity in conjunction with active learning [Jain et al., 2023b, Hernandez-Garcia et al., 2023], but also in combinatorial optimization [Zhang et al., 2023a, b], causal discovery [Deleu et al., 2022, 2023, Atanackovic et al., 2023], and probabilistic inference in general [Zhang et al., 2022b, Hu et al., 2023, 2024, Falet et al., 2024]. Although they were framed differently, GFlowNets are also deeply connected to the literature on variational inference [Malkin et al., 2023, Zimmermann et al., 2022].
A number of works have recently established connections between GFlowNets and (maximum entropy) RL. Closest to our work, Tiapkin et al. [2024] showed how the reward in MaxEnt RL can be corrected based on some backward transition probability to be equivalent to GFlowNets. Although their analysis is limited to the case where the reward function in the original soft MDP is sparse (i.e., the reward is only obtained at the end of the trajectory), they were the first to propose a correction applied at each intermediate transition as in Section 3.2. We generalized this in Theorem 3.1 with a correction at the level of the trajectories, which offers more flexibility in how the correction is distributed along the trajectory and allows intermediate rewards. Tiapkin et al. [2024] also showed similarities between GFlowNet objectives and MaxEnt RL algorithms, namely between Detailed Balance & Dueling Soft Q-Learning [Wang et al., 2016], and between Trajectory Balance & Policy Gradient [Schulman et al., 2017]. The correspondence between TB and Policy Gradient was further expanded in [Niu et al., 2024], as a direct consequence of the connections between GFlowNets, variational inference [Malkin et al., 2023], and reinforcement learning [Weber et al., 2015]. Finally, Mohammadpour et al. [2024] also introduced a correction that depends on $n(s)$ the number of (partial) trajectories to a certain state $s$, which can be learned by solving a second MaxEnt RL problem LABEL:eq:maxent-rl-problem on an “inverse” MDP. This correction corresponds to a particular choice of backward transition probability $P_{B}(s_{t}\mid s_{t+1})=n(s_{t})/n(s_{t+1})$ in Equation 8, which has the remarkable property to maximize the flow entropy.
@@ 5 Experimental results
We verify empirically the equivalences established in Section 3 on three domains: the inference over discrete factor graphs [Buesing et al., 2020], Bayesian structure learning of Bayesian networks [Deleu et al., 2022], and the generation of parsimonious phylogenetic trees [Zhou et al., 2024]. In addition to Detailed Balance (and possibly its modified version) and Trajectory Balance on the one hand (GFlowNets), and Soft Q-Learning (possibly parametrized by a policy; see Section 3.3) and Path Consistency Learning on the other hand (MaxEnt RL), we also consider a discrete version of Soft Actor-Critic [Christodoulou, 2019], which has no natural conterpart in the GFlowNet literature. For all MaxEnt RL methods, we adjust the MDP to include the correction of the reward. Note that in all the domains considered here, non-trivial intermediate rewards are available in the original MDP, meaning in particular that all instances of DB actually use the Forward-Looking formulation [Pan et al., 2023a]; the form of these intermediate rewards along with additional experimental details are available in Appendix C.